OpenStudy (anonymous):

Solve the system of equations. 2a – b + c = -5 a – b = -2 2a + b = 5

OpenStudy (anonymous):

a=1 b=3 c=-4

OpenStudy (anonymous):

frm second equation a=b-2 place the value of a in 2a+b=5 get the value of b as 3 and then a=b-2 frm 2nd so a=3-2 now place the values of a n b in equation 1 and hence get the value of c finally :p

OpenStudy (anonymous):

Thank you soo much!

OpenStudy (anonymous):

ur welcome :)